OpenStudy (anonymous):

What is the solution of 1/2x-8=-3/2x+2?

OpenStudy (anonymous):

\[1/2 x - 8 = -3/2 x + 2\]\[(x - 16)/2 = (-3x + 4)/2\] After cross multiplication, \[2x - 32 = -6x + 8\]\[2x + 6x = 8 + 32\]\[8x = 40\]\[x= 40/8\]Or\[x=5.\]

OpenStudy (anonymous):

Do you have any further queries?

OpenStudy (anonymous):

aahhh thank you! :)

OpenStudy (anonymous):

You are most welcome sis..

OpenStudy (anonymous):

@prince of persia how did u got (x-16)/2 .. if u've multiply by z denominator..2x u r supoosed to get (1-16x)/2

OpenStudy (anonymous):

@coolaidd z answer is not 5 but 1/3. firstly just multiply z equations by z denominator.. u'll get \[1-16x =-3+4x\] simply u'll get \[-12x =-4\] and z answer is 1/3

OpenStudy (anonymous):

Lgba, please check whether my answer is correct.

OpenStudy (anonymous):

soo which is it?

OpenStudy (anonymous):

As far as I know, my answer is correct. Let the critics decide it.

OpenStudy (anonymous):

ok coolaidd will know afterwards which answer was good...

OpenStudy (mimi_x3):

\[\frac{1}{2x-8} =\frac{-3}{2x+2} \] \[2x+2 = -3(2x-8)\] \[2x+2=-6x+24\] \[8x=22\] \[x = 22/8\]

OpenStudy (anonymous):

i will let you all know which one is right.. thank you!

OpenStudy (mimi_x3):

So \[x = \frac{11}{4} \] http://www.wolframalpha.com/input/?i=1%2F%282x-8%29%3D%28-3%2F%282x%2B2%29 So no one was correct I believe.

OpenStudy (anonymous):

The question was not quite clear. There should be some braces to distingiush.

OpenStudy (anonymous):

maybe u r right!

OpenStudy (anonymous):

For that reason, we all got different answers.